# NPHardEval4V: A Dynamic Reasoning Benchmark of Multimodal Large Language Models

### v2

@@ 5.2 Impact of Vision and Text Input
In evaluating the effect of different combinations of visual and textual inputs on the reasoning abilities of MLLMs, we observe varied responses across the models. Figure 9 indicates that a significant majority of the models, including both close source models like GPT-4V and open source models such as LLaVa, exhibit the highest levels of reasoning accuracy with the Default Setup, which pairs figures with limited instructional text (figure+limited_text). This suggests that their textual understanding is relatively weaker compared to their visual understanding, as the use of images to represent the question seems to be more effective than textual representation. Furthermore, when combining figure with full text, the models’ performance does not improve and instead decreases, indicating that the addition of excessive textual information may hinder their reasoning abilities. These findings highlight the importance of considering the appropriate combination of visual and textual inputs in designing tasks for MLLMs, as well as the need for further research to enhance their textual understanding abilities.
@@
In contrast, the Gemini model stands out distinctly by demonstrating superior reasoning performance with both the Text-only Setup (full_text_only) and the Vision-rich-text Setup (figure+full_text). This unique pattern indicates that Gemini may have a more advanced approach to processing and integrating comprehensive textual information, which is further enhanced by the presence of visual data. It’s noteworthy that while the addition of full textual context to the visual prompts (Vision-rich-text Setup) improves Gemini’s performance, it does not have the same positive effect on the other models tested, potentially due to the increased task complexity and vision recognition demands.
@@
These findings reveal that the interplay between visual and textual inputs is not uniform across MLLMs and that certain models may be specially attuned to leverage text, visuals, or a combination of both to maximize their reasoning proficiency. It underscores the importance of prompt design in MLLM performance and calls for further investigation into how different models process multimodal information.
6 Conclusion and Discussion
@@
In this paper, we have expanded upon the initial introduction of NPHardEval4V, a dynamic and comprehensive benchmark that scrutinizes the reasoning capabilities of Multimodal Large Language Models (MLLMs) against the backdrop of computational complexity. Our aim has been to dissect and understand the multifarious abilities of MLLMs, particularly focusing on their capacity for reasoning in response to visual-textual prompts.
@@
The experimental results, complemented by the transformed data on the impact of vision and text input, suggest that the performance of MLLMs is contingent upon not only the complexity of the task but also the nature of input they are provided. We have observed that while most models, including close source models like GPT-4V and open source ones like LLaVa, exhibit optimal performance with the Default Setup (figure with limited instructional text), it is the Gemini model that stands out, showing a marked improvement in reasoning ability when provided with text-only and vision-rich-text (figure with full textual descriptions) prompts.
@@
NPHardEval4V operates under the principle that benchmarks must be as dynamic as the models they assess. This is reflected in the benchmark’s diverse suite of tasks, which compel MLLMs to demonstrate not just recognition but sophisticated reasoning and the ability to learn adaptively. This benchmark highlights the critical need for MLLMs to process and learn from both visual and textual data effectively, underscoring the importance of dynamic evaluation tools that can truely evaluate MLLMs’ progression.
@@
In conclusion, the insights garnered from NPHardEval4V shed light on the present competencies and constraints of MLLMs. While we have detailed the specific outcomes of our experiments, the broader trends emphasize the necessity of dynamic and stringent testing to deepen our comprehension and further the advancement of AI. As we continue to extend the possibilities of MLLMs, benchmarks like NPHardEval4V will be instrumental in steering the evolution of models that are not just powerful but also multifaceted, adaptable, and truly intelligent. This study acts as a clarion call for the AI community to recognize the complexities of multimodal reasoning and to strive for models that can seamlessly integrate diverse inputs to reason and learn more like humans do.
6.1 Limitations
Inherent Bias in Model Comparisons
@@ 6.1 Limitations
The benchmarking process inherently favors models like Gemini over others due to the varied architectural strengths, which may not be fully captured in a uniform evaluation setup. Despite our attempts to level the playing field, certain models are naturally more adept at specific types of reasoning tasks due to their difference in training data, which could skew the overall comparative analysis. For example, Gemini’s superior performance in text-only setups may reflect an underlying bias in the benchmark towards linguistic reasoning over multimodal reasoning.
Limited Scope of Reasoning Tasks
Our benchmark only scratches the surface of the vast landscape of reasoning. The tasks chosen, while diverse, cannot possibly encompass the full spectrum of reasoning abilities that MLLMs might possess. There is a risk that the benchmark may overrepresent certain reasoning styles while underrepresenting others, leading to an incomplete picture of models’ true reasoning capacities.
Prompt Dependence
@@
The weaker performance of MLLMs in comparison to LLMs suggests that the integration of multimodal data may not always be beneficial for reasoning tasks. This observation implies limitations in how current models process and integrate multimodal information, potentially indicating a misalignment between model architecture and task requirements.
6.2 Research Outlook
Longitudinal Learning Studies
@@ 6.2 Research Outlook
Investigating models’ learning curves over extended periods could offer valuable insights into their potential for growth and adaptation. Long-term studies that track how models evolve with additional training data or through incremental learning can provide a deeper understanding of the underlying mechanisms that drive progress in MLLMs and identify the key factors that contribute to long-term success.
Expanding Reasoning Taxonomies
Research should continue to expand the taxonomy of reasoning tasks within benchmarks to ensure a comprehensive evaluation of MLLMs. This includes not only increasing the variety of tasks but also ensuring that they are representative of the diverse ways in which reasoning is applied in real-world contexts. Developing benchmarks that can simulate more complex, real-life reasoning scenarios will be crucial for advancing the field.
Harmonizing Model Evolution with Benchmarks
Balancing the dynamic nature of benchmarks with the pace of model development is essential. Researchers could explore introducing phased or tiered updates to benchmarks that allow for more strategic model improvements, rather than continuous, potentially disjointed, monthly updates. This approach could better align the benchmark evolution with the natural R&D cycles in AI development.
Appendix A Author Information
Author Affiliations: L. Fan, L. Li, H. Ling, J. Chi: School of Information, University of Michigan, Ann Arbor, MI 48103, US; W. Hua, M. Jin, Y. Zhang: Department of Computer Science, Rutgers University, New Brunswick, NJ 08854, US; X. Li, X. Ma: School of Control Science and Engineering, Shandong University, Jinan PRC 250061; K. Zhu, J. Wang: Microsoft Research Asia, Beijing PRC 100080.
Author Emails: lizhouf@umich.edu, wenyue.hua@rutgers.edu, lixiang0814@mail.sdu.edu.cn, v-zhukaijie@microsoft.com, u9o2n2@u.northwestern.edu, lingyaol@umich.edu, hyfrankl@umich.edu, chijk@umich.edu, jindong.wang@microsoft.com, maxin@sdu.edu.cn, yongfeng.zhang@rutgers.edu
Appendix B Recognition Accuracy

### v3

@@ 5.2. Impact of Vision and Text Input
In evaluating the effect of different combinations of visual and textual inputs on the reasoning abilities of LVLMs, we observe varied responses across the models. Figure 8 indicates that a significant majority of the models, including both close source models like Claude-3.7 and open source models such as internvl, exhibit the highest levels of reasoning accuracy with the Default Setup, which pairs figures with limited instructional text (figure+limited_text). This suggests that their textual understanding is relatively weaker compared to their visual understanding, as the use of images to represent the question seems to be more effective than textual representation. Furthermore, when combining figure with full text, the models’ performance does not improve and instead decreases, indicating that the addition of excessive textual information may hinder their reasoning abilities. These findings highlight the importance of considering the appropriate combination of visual and textual inputs in designing tasks for LVLMs, as well as the need for further research to enhance their textual understanding abilities.
@@
In contrast, the Claude-3.7 model stands out distinctly by demonstrating superior reasoning performance with both the Text-only Setup (full_text_only) and the Vision-rich-text Setup (figure+full_text). This unique pattern indicates that Claude-3.7 may have a more advanced approach to processing and integrating comprehensive textual information, which is further enhanced by the presence of visual data. It’s noteworthy that while the addition of full textual context to the visual prompts (Vision-rich-text Setup) improves Claude-3.7’s performance, it does not have the same positive effect on the other models tested, potentially due to the increased task complexity and vision recognition demands.
@@
These findings reveal that the interplay between visual and textual inputs is not uniform across LVLMs and that certain models may be specially attuned to leverage text, visuals, or a combination of both to maximize their reasoning proficiency. It underscores the importance of prompt design in MLLM performance and calls for further investigation into how different models process multimodal information.
5.3. Failed Case Analysis
While quantitative metrics show high failure rates on complex tasks, a qualitative analysis is crucial for understanding why models fail. We analyzed typical error responses and identified several common failure patterns: reasoning chain errors, instruction-parsing failures, and visual grounding errors.
Below, we present a sentence-by-sentence analysis of a typical erroneous response for a Knapsack Problem (KSP) instance.
1Task: Given 9 items with weights {9, 6, 1, 1, 7, 6, 4, 9, 6} and values {8, 8, 4, 6, 7, 8, 6, 8, 1}, select a subset of items to maximize total value without exceeding a knapsack capacity of 45.
2Correct Answer: Select items with weights {9, 6, 1, 1, 7, 6, 4, 9} (items 0, 1, 2, 3, 4, 5, 6, 7) for a total value of 55.
3Typical Incorrect Model Response:
4<root>
5<reasoning>
[…]
28The optimal subset is {0, 1, 2, 3, 4, 5, 6} summing up to weight 45 and total value 47.
29 </reasoning> <final_answer>{’SelectedItems’: [0,1,2,3,4,5,6], TotalValue’: 47}</final_answer> </root>
5.3.1. Correct Intermediate Steps, Wrong Reasoning Strategy:
The model correctly identifies all items, their weights, and values from the input. It also parses the knapsack capacity and performs valid arithmetic to ensure the selected items’ total weight does not exceed the constraint (the sum is exactly 45). Furthermore, the computed total value for the selected items—47 for items 0, 1, 2, 3, 4, 5, 6—is accurate. However, while these intermediate calculations are correct, the overall selection strategy is suboptimal. The model’s method of sorting items by value/weight (“efficiency”) and then applying a dynamic programming approach misses the globally optimal solution.
5.3.2. Source of Failure:
Reasoning Chain Error. The primary source of failure is in the reasoning strategy. Although the model claims to use dynamic programming, the actual logic reflects a greedy or near-greedy selection process based on item efficiency. This heuristic does not guarantee an optimal solution for the Knapsack problem. In this example, the model selects items 0, 1, 2, 3, 4, 5, 6 (total value 47), but fails to include item 7, which would increase the total value to 55 without exceeding capacity. The model does not properly explore all combinations to maximize the objective, demonstrating an incomplete implementation of dynamic programming or a misunderstanding of the required exhaustive search for optimality.
5.3.3. Discussion of Hallucination:
This example illustrates a form of hallucination: the model outputs a reasoning trace that appears algorithmic and plausible, but actually misrepresents both its method (“using dynamic programming”) and its effectiveness. The step-by-step explanation is internally consistent but ultimately implements a shortcut that cannot always solve the problem optimally. This highlights the risk of accepting well-written explanations at face value—surface plausibility may mask underlying errors in logic or algorithm selection.
6. Conclusion and Discussion
@@
In this paper, we have expanded upon the initial introduction of NPHardEval4V, a dynamic and comprehensive benchmark that scrutinizes the reasoning capabilities of Large Vision-Language Models (LVLMs) against the backdrop of computational complexity. Our aim has been to dissect and understand the multifarious abilities of LVLMs, particularly focusing on their capacity for reasoning in response to visual-textual prompts.
@@
The experimental results, complemented by the transformed data on the impact of vision and text input, suggest that the performance of LVLMs is contingent upon not only the complexity of the task but also the nature of input they are provided. We have observed that while most models, including close source models like Claude-3.7 and open source ones like Internvl, exhibit optimal performance with the Default Setup (figure with limited instructional text), it is the Gemini model that stands out, showing a marked improvement in reasoning ability when provided with text-only and vision-rich-text (figure with full textual descriptions) prompts.
@@
NPHardEval4V operates under the principle that benchmarks must be as dynamic as the models they assess. This is reflected in the benchmark’s diverse suite of tasks, which compel LVLMs to demonstrate not just recognition but sophisticated reasoning and the ability to learn adaptively. This benchmark highlights the critical need for LVLMs to process and learn from both visual and textual data effectively, underscoring the importance of dynamic evaluation tools that can truely evaluate LVLMs’ progression.
@@
In conclusion, the insights garnered from NPHardEval4V shed light on the present competencies and constraints of LVLMs. While we have detailed the specific outcomes of our experiments, the broader trends emphasize the necessity of dynamic and stringent testing to deepen our comprehension and further the advancement of AI. As we continue to extend the possibilities of LVLMs, benchmarks like NPHardEval4V will be instrumental in steering the evolution of models that are not just powerful but also multifaceted, adaptable, and truly intelligent. This study acts as a clarion call for the AI community to recognize the complexities of multimodal reasoning and to strive for models that can seamlessly integrate diverse inputs to reason and learn more like humans do.
6.1. Limitations
@@
The benchmarking process inherently favors models like claude-3.7 over others due to the varied architectural strengths, which may not be fully captured in a uniform evaluation setup. Despite our attempts to level the playing field, certain models are naturally more adept at specific types of reasoning tasks due to their difference in training data, which could skew the overall comparative analysis. For example, claude-3.7’s superior performance in text-only setups may reflect an underlying bias in the benchmark towards linguistic reasoning over multimodal reasoning.
Furthermore, model selection is concentrated on mainstream large models, lacking the inclusion of models that leverage specialized visual priors, such as SAM or DETR-ViT based architectures. Evaluating such models could reveal different performance patterns. Similarly, the diversity of tasks on the benchmark is limited. The current tasks are grounded in classic algorithmic problems and do not cover other important areas like image semantic reasoning, visual common-sense reasoning, chart understanding, or multi-step procedure generation from visual guides. This may lead to an incomplete picture of models’ true reasoning capacities.
The design of image and text inputs, despite our efforts to ensure informational equivalence, may still introduce biases. Images can simplify certain details, which makes some models ’visually easier to answer.’ To properly calibrate this effect, future work should add a human baseline or a o3 + OCR baseline. This would help determine whether humans or other powerful systems also exhibit a preference for a particular modality, providing a stronger reference for assessing fairness.
Finally, evaluation metrics focus on accuracy-based indicators (RA, ER, AA). While useful, they do not provide a deep view into the model’s internal processes. The analysis lacks a systematic evaluation of the interpretability of reasoning chains, which could be enhanced by visualizing the model’s reasoning trace or performing attribution analysis.
@@
The weaker performance of LVLMs in comparison to LLMs suggests that the integration of multimodal data is not always be beneficial for reasoning tasks. This observation implies limitations in how current models process and integrate multimodal information, indicating a misalignment between model architecture and task requirements.
6.2. Research Outlook
@@
Investigating models’ learning curves over extended periods could offer valuable insights into their potential for growth and adaptation. Long-term studies that track how models evolve with additional training data or through incremental learning can provide a deeper understanding of the underlying mechanisms that drive progress in LVLMs and identify the key factors that contribute to long-term success.
To complement long-term studies of model development, future work should also broaden the taxonomy of reasoning tasks. This includes incorporating challenges such as visual common-sense reasoning, chart and diagram interpretation, and multi-step procedural generation—areas that better reflect the diverse applications of modern LVLMs.
Alongside task expansion, deeper interpretability is crucial. One direction is to introduce reasoning trace visualizations, enabling researchers to inspect the model’s step-by-step logic and identify points of failure. Another is to strengthen attribution analysis—particularly of the vision module—using techniques such as image masking or attention map inspection to reveal which visual elements the model relies on during reasoning.
These efforts can enrich the understanding of model behavior and better inform the design of benchmarks. As such, benchmark evolution should be carefully paced. Rather than frequent, disjointed updates, a phased or tiered update cycle could synchronize better with the natural R&D rhythm, encouraging more strategic improvements in model design and evaluation.
Appendix A Details of Reasoning Tasks
There are nine tasks in total in our benchmark and each complexity class have three unique problem categories.
A.1. P (Polynomial time) Tasks
This class consists of tasks that can be solved by a deterministic Turing machine in polynomial time. Essentially, it represents tasks that are efficiently solvable. We include three P problems in the benchmark, namely Binary Search Problem (BSP), Edit Distance Problem (EDP), and Shortest Path Problem (SPP).
A.1.1. The Binary Search Problem (BSP)
BSP is defined by the task of locating the index of a target value within a given array as if it were sorted. The problem is presented with an unsorted array $A$ of $n$ elements and a target value $T$. The objective is to determine the index $i$ that $T$ would occupy if $A$ were sorted in ascending order. Formally, after sorting $A$, the goal is to find an index $i$ such that $A[i]=T$, or to conclude that the target is not present. This process inherently involves two stages: sorting the array and then searching for the target. The search phase is efficiently handled by algorithms like binary search, which iteratively reduces the search space and operates in logarithmic time. BSP is structured into 10 complexity levels. These levels scale by increasing the array length by one (from 3 to 12) and expanding the upper limit of the number range by 5 (from (1, 15) to (1, 60)) at each level.
A.1.2. Edit Distance Problem (EDP)
EDP is about finding the minimum number of operations required to transform one string into another. Given two strings, $A$ and $B$, of lengths $m$ and $n$ respectively, the aim is to determine the minimum number of operations needed to convert $A$ into $B$. The allowable operations are insertion, deletion, and substitution of a single character. Formally, the problem can be defined as finding a minimum number $d$ such that string $A$ can be transformed into string $B$ using $d$ operations. This algorithm has a time complexity of $\mathcal{O}(ab)$ where $a$ and $b$ are the lengths of the strings. When the full dynamic programming table is constructed, its space complexity is also $\mathcal{O}(ab)$. EDP has widespread applications, especially in fields like computational biology for sequence alignment, natural language processing for spell checking and correction, and in data analysis for measuring similarity between data strings. EDP is structured into 10 complexity levels, tailored to measure the minimal number of edits required to transform one string into another. Each level is characterized by two strings whose lengths are equal and progressively increase from 3 to 12 characters from Level 1 to Level 10. Concurrently, the character range for constructing these strings is expanded, starting with the first 6 letters and extending by 2 additional letters at each subsequent level.
A.1.3. Shortest Path Problem (SPP)
SPP is about finding the shortest path between two nodes in a non-negative weighted graph. In our experiments, we ask for the shortest path between the first and last nodes. Given a graph $G=(V,E)$ with a weight function $w:E\rightarrow\mathbb{R}$ assigning weights to edges, and two vertices $u$ and $v$ in $V$, the task is to find the path from $u$ to $v$ that minimizes the total weight. This is often solved using Dijkstra’s algorithm which systematically expands the shortest path from the starting node until it reaches the target node. Formally, the problem is to find a path $P=(v_{1},v_{2},...,v_{k})$, where $v_{1}=u$ and $v_{k}=v$, such that the sum of weights of consecutive edges in $P$, $\sum_{i=1}^{k-1}w(v_{i},v_{i+1})$, is minimized. This problem can be used in network routing, GPS navigation systems, and logistics to find the shortest or most efficient path between two points. It helps in reducing travel time and costs in transportation and communication networks. SPP involves determining the shortest path across different complexity levels, ranging from Level 1 to Level 10. Each level is characterized by an increasing number of nodes (starting from 4 to 13), edges (5 to 14), and a maximum weight (6 to 15) that escalates linearly with each level. The complexity of the problem scales by adding one node, one edge, and increasing the maximum weight by one for each subsequent level.
A.2. NP-complete problems
This is a subset of NP. A problem is NP-complete if it is in NP and as hard as any problem in NP. If any NP-complete problem can be solved in polynomial time, then every problem in NP can also be solved in polynomial time. We include three NP-complete problems that are not in P in the benchmark, namely Traveling Salesman Problem Decision Version (TSP-D), Graph Coloring Problem Decision Version (GCP-D), and Knapsack Problem (KSP).
A.2.1. Traveling Salesman Problem (Decision Version, TSP-D)
TSP-D is concerned with determining if a salesman can complete a route, visiting each city at least once, with the total travel distance being less than a specified value. Given a complete graph $G=(V,E)$ with vertices $V$ representing cities and edges $E$ representing paths between cities, each edge $(i,j)$ is assigned a distance $d(i,j)$. The decision version of this problem asks whether there exists a tour (a sequence of cities) such that the total distance of the tour is less than or equal to a given value $D$. Formally, the problem can be stated as finding a permutation $P$ of the set of cities 1,2,,n{1,2,...,n} that satisfies the condition $\sum_{i=1}^{n-1}d(P(i),P(i+1))+d(P(n),P(1))\leq D$. This problem is useful in logistics and supply chain management in planning efficient delivery routes and schedules (Roberti and Ruthmair, 2021). TSP-D configuration spans 10 complexity levels with node counts from 4 to 13, similar to the TSP. It also introduces a threshold of 0.75, setting factor of the allowed travel distance to the total possible distance.
A.2.2. Graph Coloring Problem (Decision Version, GCP-D)
GCP-D involves determining if it is possible to color the vertices of a graph using a given number of colors so that no two adjacent vertices share the same color. Given an undirected graph $G=(V,E)$, with $V$ representing vertices and $E$ representing edges, the goal is to find out if there is a way to assign one of $k$ colors to each vertex such that for any edge $(u,v)\in E$, the vertices $u$ and $v$ have different colors. The formal statement is to determine if there exists a coloring function $c:V\to{1,2,...,k}$ such that for every edge $(u,v)\in E$, $c(u)\neq c(v)$. It has wide applications in Round-Robin Sports Scheduling, Aircraft scheduling, and Biprocessor tasks (Ahmed, 2012). GCP-D has difficulty levels 1 to 10 with questions of 6, 8, 10, 12, 14, 16, 18, 20, 22, and 24 average edges and 6, 7, 8, 9, 10, 11, 12, 13, 14, and 15 nodes. Beginning with graphs of 6 nodes and 6 edges, each subsequent level incorporates an additional 2 edges and 1 node, culminating in graphs of 24 edges and 15 nodes.
A.2.3. Knapsack Problem (KSP)
KSP asks whether a subset of items can be chosen to fit into a knapsack of fixed capacity without exceeding it, while also maximizing the total value of the selected items. Consider a set of items, each with a weight $w_{i}$ and a value $v_{i}$, and a knapsack with a weight capacity $W$. The problem is to select a subset of these items such that the total weight does not exceed $W$ and the total value is maximized. Formally, let $x_{i}$ be a binary variable indicating whether item $i$ is included in the knapsack ($x_{i}=1$) or not ($x_{i}=0$). The problem can be stated as maximizing $\sum_{i=1}^{n}v_{i}x_{i}$ subject to the constraint $\sum_{i=1}^{n}w_{i}x_{i}\leq W$, where $n$ is the number of items. It is used in resource allocation and budgeting where the goal is to maximize the total value of a selection under a weight or cost constraint. Applications include cargo loading, and electric vehicle charging (Sun et al., 2020; Cho, 2019). KSP is organized into 10 levels, each marked by an ascending number of items (from 4 to 13), a weight range (1 to the level number), a value range (identical to the weight range), and a knapsack capacity that starts at 20 and increases by 5 with each level. The complexity elevates by broadening the weight and value ranges, adding more items, and enlarging the knapsack’s capacity.
A.3. NP-hard problems
These problems are at least as hard as the hardest problems in NP. They may not necessarily be in NP (i.e., they may not have solutions verifiable in polynomial time) but solving an NP-hard problem in polynomial time would imply that P = NP. We include three NP-hard problems that are not reducible to NP-complete problems in the benchmark, namely Traveling Salesman Problem Optimization Version (TSP), Graph Coloring Problem Optimization Version (GCP), and Meeting Scheduling Problem (MSP).
A.3.1. Traveling Salesman Problem (Optimization Version, TSP)
TSP-O involves finding the shortest route for a salesman to visit each city exactly once and return to the starting city. Given a complete graph $K_{n}$ with $n$ vertices, where each vertex represents a city and each edge $(i,j)$ is assigned a non-negative cost or distance $d(i,j)$, the problem is to find the shortest possible route that visits each city exactly once and returns to the origin city. Formally, let $P$ be a permutation of the set of cities 1,2,,n{1,2,...,n} representing the order in which the cities are visited. The traveling salesman problem can be formulated as finding the permutation $P$ that minimizes the total travel cost, given by the function $f(P)=d(P(n),P(1))+\sum_{i=1}^{n-1}d(P(i),P(i+1))$. This problem is important in operational research and logistics to find the most efficient route to visit multiple locations and return to the origin, particularly route planning for delivery services, maintenance operations, and sales. TSP is structured across 10 complexity levels, each defined by a set number of nodes ranging from 4 to 13. The complexity of the problem increases with the addition of more nodes, enhancing the challenge of finding the shortest possible route that visits each node exactly once and returns to the starting point.
A.3.2. Graph Coloring Problem (Optimization Version, GCP)
GCP-O refers to the problem of coloring vertices of a graph in such a way that no two adjacent vertices have the same color. Given an undirected graph $G=(V,E)$, where $V$ is the set of vertices and $E$ is the set of edges, assign a color to each vertex such that no two adjacent vertices have the same color. Formally, let $c:V\to C$ be a function that assigns a color from a set of colors $C$ to each vertex in $V$. The graph coloring problem can be formulated as finding a proper coloring, i.e., a function $c$ such that for every edge $(u,v)\in E$, $c(u)\neq c(v)$. This problem is used in constraint satisfaction problems and applied in exam timetabling and register allocation in compilers (Lintzmayer et al., 2011). GCP is devised with 10 levels of complexity, featuring questions with average edges ranging from 6 to 24 and nodes from 6 to 15. Beginning with graphs of 6 nodes and 6 edges, each subsequent level incorporates an additional 2 edges and 1 node, culminating in graphs of 15 nodes and 24 edges, progressively elevating the difficulty of assigning colors to each node without any two adjacent nodes sharing the same color.
A.3.3. Meeting Scheduling Problem (MSP)
MSP deals with allocating time slots for meetings such that all constraints, including participant availability and room capacity, are satisfied without overlaps. Given a set of $n$ participants and their availability for $m$ time slots, find a schedule that maximizes the number of participants who can attend the meeting. Formally, let $A={a_{1},a_{2},...,a_{n}}$ be the set of participants and $T={t_{1},t_{2},...,t_{m}}$ be the set of time slots. For each participant $a_{i}$, let $S_{i}$ be a subset of $T$ representing the times when $a_{i}$ is available and $m_{i}$ be a subset of meetings that are required to attend. The meeting scheduling problem can be formulated as finding a subset $S\subseteq T$ such that $|{a_{i}\in A|S_{i}\cap S\neq\emptyset}|$ is maximized. In other words, the aim is to find a scheduling subset $S_{i}$ where the collective availability of participants intersects with $S_{i}$, ensuring maximum participation. This problem is crucial in organizational management for scheduling meetings involving multiple participants with varying availability. It ensures optimal utilization of time and resources and is used in corporate scheduling systems and collaborative software (Bofill et al., 2022). MSP outlines complexity across 10 levels, determined by an increasing number of meetings (2 to 11), participants (one more than the number of meetings, i.e., 3 to 12), and time slots (two more than the number of meetings, i.e., 4 to 13). Each level advances the problem’s complexity by adding one meeting, consequently increasing the number of participants and time slots required for scheduling.
Appendix B NPHardEval Benchmark Consistency
The NPHardEval Benchmark is designed for high consistency, a feature verified through preliminary assessments and supported by ongoing developments for open-source checks. For each reasoning task, such as the Binary Search Problem (BSP), the benchmark comprises three distinct versions (V0, V1, and V2) to assess the stability of a model’s problem-solving capabilities. We tested this on twelve open-source models — DeepSeekVL2, Intern, Qwen2, Ovis, Llava_Qwen, NVLM, PointsTest, Ola_7B, Llama90BVision, GLM, Qwen25VL72B, and Gemma — as well as proprietary models including GPT-4o-2024-11-20, Gemini 2.0 Flash Exp, and Claude 3.7 Sonnet-20250219. By calculating the performance variance on each problem type across the benchmark versions, we generated 135 data points for model-specific consistency analysis. The summary statistics are as follows:
  • Minimum variance: 0.0 (approximately half of the variances are near zero)
  • Mean variance: 0.01167
  • Maximum variance: 0.11029
These statistics indicate minimal variance across the benchmark versions, suggesting a high level of consistency with each update.
Appendix C Recognition Accuracy